OpenStudy (anonymous):

Find the zeros in the function f(x)= 3x^2 +9x +6/ 3x-3

OpenStudy (anonymous):

Start by factoring the top

OpenStudy (anonymous):

Thats the only problem i'm having i cant find the numbers to factor it

OpenStudy (anonymous):

i tried (3x-3)(x-2) that doesnt work

OpenStudy (anonymous):

I get 3(x+1)(x+2) when I factor it

OpenStudy (anonymous):

After you factor the bottom you get 3(x-1)

OpenStudy (anonymous):

So then set your parenthesis to zero to find the zeros x+1=0 x=-1 x+2=0 x=-2 x-1=0 x+1 So it appears the zeros are 1, -1, and -2